OpenStudy (ksaimouli):

3x^2+3x+2y=0 complete the square and reduce to one standard form y-b=A(x-a)^2 or x-a=A(y-b)^2

OpenStudy (callisto):

\[3x^2+3x+2y=0\] \[3(x^2+x)=-2y\] \[3[x^2+x + (\frac{1}{2})^2-(\frac{1}{2})^2]=-2y\]\[3[x^2+x + (\frac{1}{2})^2-(\frac{1}{2})^2]=-2y\]\[3(x+\frac{1}{2})^2-\frac{3}{4}=-2y\]Divide both sides by -2\[-\frac{3}{2}(x+\frac{1}{2})^2+\frac{3}{8}=y\]\[-\frac{3}{2}(x+\frac{1}{2})^2=y-\frac{3}{8}\]\[y-\frac{3}{8}=-\frac{3}{2}(x+\frac{1}{2})^2\]

OpenStudy (ksaimouli):

why did u squared 1/2

OpenStudy (ksaimouli):

i think u should add 1/2 on both sides

OpenStudy (callisto):

Because of the identity \((a+b)^2 = a^2+2ab+b^2\) For the expression x^2 + x , a=x, 2ab = x So, 2xb = x b = 1/2 To complete square, we need to add b^2, that is (1/2)^2 = 1/4 But we can't just add 1/4. So, I add it and subtract it, like this: \[3[x^2+x + (\frac{1}{2})^2-(\frac{1}{2})^2]=-2y\] Actually, adding a term to both sides is EQUAL to adding the term and subtracting the same term on the SAME side. Like this: a = 5 , assume we add b to both sides a + b = 5 + b a + b - b = 5 See ?

OpenStudy (ksaimouli):

and last question how did u get (1/2)^2 = 3/4

OpenStudy (ksaimouli):

IN 4TH STEP

OpenStudy (callisto):

There are 2 step to get that value. \[3[x^2+x + (\frac{1}{2})^2-(\frac{1}{2})^2]=-2y\]\[3(x^2+x + (\frac{1}{2})^2)-3(\frac{1}{2})^2=-2y\]\[3(x+\frac{1}{2})^2-3(\frac{1}{4})=-2y\]\[3(x+\frac{1}{2})^2-(\frac{3}{4})=-2y\]

OpenStudy (ksaimouli):

where did u get that -3

OpenStudy (callisto):

Distributive.

OpenStudy (ksaimouli):

hoooo u got got it from the first 3